OpenStudy (anonymous):

1-cos x=sin x in the range of [0,2pi)

OpenStudy (anonymous):

cosx+sinx=1 then (cosx+sinx)^2=1

OpenStudy (anonymous):

can you tell what is the next ?

OpenStudy (anonymous):

sin x = ( 1- cos ^2 x) ^(1/2) ( 1- cos ^2 x) ^(1/2) = 1+cos x

OpenStudy (anonymous):

(cosx+sinx)^2=1=cos^2(x)+sin^2(x)+2sinx cosx=1 and it is a rule cos^2?(theta)+sin^2(theta)=1 so (cos^2x+sin^2x+2sinxcosx=1+sin2x=1 then sin2x=0

OpenStudy (anonymous):

this pic shows the graph of sin2x u see sin2x=0 in the period (0,2pi) at x=90,180,270,360 degrees

OpenStudy (anonymous):

and x=0 also